A 74.2 kg skier is at rest at the top of a 268 m high run. The slope is inclined at a 24.4º angle relative to horizontal. How much kinetic energy does the skier have?
1
Expert's answer
2020-02-04T09:22:29-0500

Kinetic energy = change in potential energy = "m\times g\times \Delta h"

Kinetic energy = 74.2 * 10 * 268 = 198856 joules